Convert the following ratios to percentages
2:3
Convert the following ratios to percentages
2:3

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

2:3=2/3 =2/3 ×100/100 = 2/3 ×100%=0.666×100% = 66.66%= 66 whole 2/3 %

Thanks